OpenStudy (anonymous):

Find a cubic function with the given zeros. Square root of seven., -Square root of seven., -4

OpenStudy (anonymous):

\[\sqrt{7}, -\sqrt{7}, -4\]

OpenStudy (phi):

you are told x= sqr(7), or x-sqr(7)= 0 x= -sqr(7) or x+ sqr(7)=0 x= -4 or x+4=0 multiply out \[ (x-\sqrt{7})(x+\sqrt{7})(x+4) \] I would first do \[(x-\sqrt{7})(x+\sqrt{7}) \]

OpenStudy (anonymous):

would that be \[x^{2}-7\] @phi

OpenStudy (anonymous):

then\[x ^{3}+4x ^{2}-7x-28\]?

OpenStudy (anonymous):

what's the question

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

@kirarittgers

OpenStudy (anonymous):

did u get it @kirarittgers

OpenStudy (anonymous):

no lol but i have the answer so thank youuu